# Gaussian Process Implicit Surfaces as Control Barrier Functions for Safe Robot Navigation

### v1

@@ VI-B Chair Implicit Surface Modeling
where $t=\frac{\sqrt{3}\ \|\ \mathbf{x}_{i}-\mathbf{x}_{j}\|}{l}\in\mathbb{R}_{\geq 0}$, $\mathbf{x}_{i},\mathbf{x}_{j}\in\mathbb{R}^{n}$, $i,j\in\{1,\dots,N\}$, $N$ is the number of samples, and $\sigma_{f}\in\mathbb{R}$, $l\in\mathbb{R}$, $\sigma_{\omega}$ are signal variance, length scale, and signal noise hyperparameters respectively. Unlike the SE kernel, which assumes infinite smoothness and may be unrealistic for physical phenomena, the Matérn class allows explicit control of smoothness and better captures real-world structures [31].
@@ VI-D Safe Control Synthesis for Quadrotor
We use thrust-attitude setpoints to control the Crazyflie [32]. We use the same second-order integrator and rectification process as [33]. The relative degree for the second-order integrator is $\rho=2$. Hence, the associated Lie derivatives for the Gaussian CBF (with and without sparsity) in (21) are,
@@
where $a_{\mathrm{gp}}^{i}$ and $a_{\mathrm{sgp}}^{i}$ are the $i^{\mathrm{th}}$ entries of $\mathbf{y}_{N}^{\top}\mathbf{\overline{K}\hskip 1.42271pt}^{-1}\hskip-5.69046pt\in\mathbb{R}^{1\times N}$ and $\hskip-5.69046pt\mathbf{y}_{N}^{\top}\big(\mathbf{Q}_{M}^{-1}\ \mathbf{K}_{MN}\big(\mathbf{\Lambda}_{N}+\sigma_{y}^{2}\mathbf{I}_{N}\big)^{-1}\big)^{\top}\hskip-5.69046pt\in\hskip-2.84544pt\mathbb{R}^{1\times M}\hskip-8.5359pt$ , respectively. $b_{\mathrm{gp}}^{i}(\mathbf{x})$ and $b_{\mathrm{sgp}}^{i}(\mathbf{x})$ are the $i^{\mathrm{th}}$ entries of $\mathbf{k}(\mathbf{x})^{\top}\mathbf{\overline{K}\hskip 1.42271pt}^{-1}\hskip-5.69046pt\in\mathbb{R}^{1\times N}$ and $\mathbf{k}_{M}(\mathbf{x})^{\top}\big(\mathbf{K}_{M}^{\ -1}-\mathbf{Q}_{M}^{\ -1}\big)\in\mathbb{R}^{1\times M}$ , respectively. $\mathbf{A}_{\mathrm{gp}}=\nabla\mathbf{k}(\mathbf{x})\mathbf{\overline{K}\hskip 2.84544pt}^{-1}\nabla\mathbf{k}(\mathbf{x})^{\hskip-1.42271pt\top}\in\mathbb{R}^{N\times N}$, $\mathbf{A}_{\mathrm{sgp}}=\hskip-2.84544pt\nabla\mathbf{k}_{M}(\mathbf{x})\big(\mathbf{K}_{M}^{-1}\hskip-4.26773pt-\hskip 0.0pt\mathbf{Q}_{M}^{-1}\big)\hskip-1.42271pt\nabla\mathbf{k}_{M}(\mathbf{x})^{\top}\in\mathbb{R}^{M\times M}$, $\nabla\mathbf{k}(\mathbf{x})=\frac{\partial\mathbf{k}(\mathbf{x})}{\partial\mathbf{x}}^{\top}\in\mathbb{R}^{n\times N}$, $\nabla\mathbf{k}_{M}(\mathbf{x})=\frac{\partial\mathbf{k}(\mathbf{x})}{\partial\mathbf{x}}^{\top}\in\mathbb{R}^{n\times M}$, and $\frac{\partial^{2}\mathbf{k}_{(i)}(\mathbf{x})}{\partial\mathbf{x}^{2}}$ is the partial derivative of (16) with respect to $\mathbf{x}$. Since the relative degree is $2$, we adopt higher-order CBFs, in particular the exponential CBF [1, 34]. Using the Lie derivatives in (VI-D), we can design a QP of the form (18) to yield rectified control inputs that guarantee forward invariance of the safe set.

### v2

@@ VI-B Chair Implicit Surface Modeling
where $t=\frac{\sqrt{3}\ \|\ \mathbf{x}_{i}-\mathbf{x}_{j}\|}{l}\in\mathbb{R}_{\geq 0}$, $\mathbf{x}_{i},\mathbf{x}_{j}\in\mathbb{R}^{n}$, $i,j\in\{1,\dots,N\}$, $N$ is the number of samples, and $\sigma_{f}\in\mathbb{R}$, $l\in\mathbb{R}$, $\sigma_{\omega}$ are signal variance, length scale, and signal noise hyperparameters respectively. Unlike the SE kernel, which assumes infinite smoothness and may be unrealistic for physical phenomena, the Matérn class allows explicit control of smoothness and better captures real-world structures [math_interpolation_Stein1999].
@@ VI-D Safe Control Synthesis for Quadrotor
We use thrust-attitude setpoints to control the Crazyflie [quadrotor_trajectory_generation_control_Mellinger2012]. We use the same second-order integrator and rectification process as [cbf_teleoperation_Xu2018]. The relative degree for the second-order integrator is $\rho=2$. Hence, the associated Lie derivatives for the Gaussian CBF (with and without sparsity) in (21) are,
@@
where $a_{\mathrm{gp}}^{i}$ and $a_{\mathrm{sgp}}^{i}$ are the $i^{\mathrm{th}}$ entries of $\mathbf{y}_{N}^{\top}\mathbf{\overline{K}\hskip 1.42271pt}^{-1}\hskip-5.69046pt\in\mathbb{R}^{1\times N}$ and $\hskip-5.69046pt\mathbf{y}_{N}^{\top}\big(\mathbf{Q}_{M}^{-1}\ \mathbf{K}_{MN}\big(\mathbf{\Lambda}_{N}+\sigma_{y}^{2}\mathbf{I}_{N}\big)^{-1}\big)^{\top}\hskip-5.69046pt\in\hskip-2.84544pt\mathbb{R}^{1\times M}\hskip-8.5359pt$ , respectively. $b_{\mathrm{gp}}^{i}(\mathbf{x})$ and $b_{\mathrm{sgp}}^{i}(\mathbf{x})$ are the $i^{\mathrm{th}}$ entries of $\mathbf{k}(\mathbf{x})^{\top}\mathbf{\overline{K}\hskip 1.42271pt}^{-1}\hskip-5.69046pt\in\mathbb{R}^{1\times N}$ and $\mathbf{k}_{M}(\mathbf{x})^{\top}\big(\mathbf{K}_{M}^{\ -1}-\mathbf{Q}_{M}^{\ -1}\big)\in\mathbb{R}^{1\times M}$ , respectively. $\mathbf{A}_{\mathrm{gp}}=\nabla\mathbf{k}(\mathbf{x})\mathbf{\overline{K}\hskip 2.84544pt}^{-1}\nabla\mathbf{k}(\mathbf{x})^{\hskip-1.42271pt\top}\in\mathbb{R}^{N\times N}$, $\mathbf{A}_{\mathrm{sgp}}=\hskip-2.84544pt\nabla\mathbf{k}_{M}(\mathbf{x})\big(\mathbf{K}_{M}^{-1}\hskip-4.26773pt-\hskip 0.0pt\mathbf{Q}_{M}^{-1}\big)\hskip-1.42271pt\nabla\mathbf{k}_{M}(\mathbf{x})^{\top}\in\mathbb{R}^{M\times M}$, $\nabla\mathbf{k}(\mathbf{x})=\frac{\partial\mathbf{k}(\mathbf{x})}{\partial\mathbf{x}}^{\top}\in\mathbb{R}^{n\times N}$, $\nabla\mathbf{k}_{M}(\mathbf{x})=\frac{\partial\mathbf{k}(\mathbf{x})}{\partial\mathbf{x}}^{\top}\in\mathbb{R}^{n\times M}$, and $\frac{\partial^{2}\mathbf{k}_{(i)}(\mathbf{x})}{\partial\mathbf{x}^{2}}$ is the partial derivative of (16) with respect to $\mathbf{x}$. Since the relative degree is $2$, we adopt higher-order CBFs, in particular the exponential CBF [cbf_theory_Ames2019control, cbf_exponential_Nguyen2016]. Using the Lie derivatives in (VI-D), we can design a QP of the form (18) to yield rectified control inputs that guarantee forward invariance of the safe set.